A sail boat is traveling east at a speed of 10 mph while the wind is blowing at 20 mph from the north.(a) At what speed and (b) in what direction does the sailor feel the relative wind blowing in his face ?
Answer #1

here,

speed of boat , v1 = 10 mph i

speed of wind , v2 = 20 mph j

the speed of wind relative to sailor , v21 = v2 - v1

v21 = - 10 mph i + 20 mph j

a)

the speed of boat , |v| = sqrt(v1^2 + v2^2)

|v| = sqrt(10^2 + 20^2) = 22.4 mph

b)

the direction , theta = arctan(20/10)

theta = 63.4 degree North of West